OpenStudy (anonymous):

Someone help please Now that you have helped Bernard design the pool, he presents you with a situation. There are x number of communities that will be designing homes with the pools. The number of homes within each community is six more than the number of communities. The number of hours that will be required to build each home is double the number of homes in each community. B. Using the three expressions, interpret the meaning of each grouping of factors as a single unit, and also simplify each expression by finding their product.

OpenStudy (anonymous):

C. Explain how polynomials demonstrated the closure property throughout this exploration. I've already solved A

OpenStudy (anonymous):

@hari5719 @SolomonZelman @zepdrix @triciaal @mathmate @mathstudent55 @mathmale @pooja195 @Awolflover1 @satellite73 @ParthKohli

OpenStudy (anonymous):

@sleepyjess @dan815 @SithsAndGiggles

OpenStudy (anonymous):

@boldjon @DannyO19

OpenStudy (boldjon):

might be long but take ur time tell me if u don't get it x = number of communities x + 6 = number of houses per community 2(x + 6) = Hours required to build Alright, so you got the above information. Now, the question asks you to combine them via multiplication to get some other measure. For example: If we take the expression "x" for the number of communities and the expression (x + 6) for the number of houses per community, we can multiply them together to yield: x(x+6) = x^2 + 6x = The number of houses altogether You can do the same for 2(x+6) and (x+6), as multiplying them together will yield the total time needed to build all the houses in a single community: 2(x+6)(x+6) = 2(x^2 + 12x + 36) = 2x^2 + 24x + 72 is the expression for the amount of hours needed per community. The question asks for a grouping of all the the above expressions. To do this, you do the same as I showed above, but you use all three expressions. This will yield the total time needed to build all houses in all of the communities: 2(x+6) * (x+6) * x = 2x^2 + 24x + 72 [from previous] (x) = 2x^3 + 24x^2 + 72x As for the closure property question, all of the products that we just generated are polynomials. So a multiplication of polynomials yields a closed set of polynomials since their multiplication yields another polynomial. This can be understood through the simple addition of exponents when multiplying by said exponents, and since non-negative exponents are present, their results must also be positive (since you can't subtract exponents under these conditions and get a non-polynomial answer).

OpenStudy (anonymous):

Alright, @boldjon I'm gonna try understanding what you just said, tell me if I'm wrong

OpenStudy (boldjon):

k

OpenStudy (boldjon):

oh one more thing SUMMARY: x = # of communities x + 6 = H 2H = T

OpenStudy (anonymous):

So the grouping thing in Part B just wants me to multiply the number of hours * the number of houses * the number of communities, correct?

OpenStudy (boldjon):

yup

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Wuts the difference between that and: and also simplify each expression by finding their product. ???

OpenStudy (anonymous):

Oh wait nvm

OpenStudy (boldjon):

XD u got it?

OpenStudy (anonymous):

Basically, I am supposed to give equations, like the number of hours per house could be calculated using: 2(x + 6) In that, I could simplify it to 2x + 12, that's what that part is asking, right?

OpenStudy (boldjon):

yeah, it's easy once u get it

OpenStudy (anonymous):

Alright. Now Part C

OpenStudy (anonymous):

Too much work scrolling up but: "As for the closure property question, all of the products that we just generated are polynomials. So a multiplication of polynomials yields a closed set of polynomials since their multiplication yields another polynomial. This can be understood through the simple addition of exponents when multiplying by said exponents, and since non-negative exponents are present, their results must also be positive (since you can't subtract exponents under these conditions and get a non-polynomial answer). "

OpenStudy (boldjon):

O_O part c?

OpenStudy (anonymous):

Yea, we only did Part B

OpenStudy (anonymous):

I'm analyzing ur block of text that answers Part C, hold on "As for the closure property question, all of the products that we just generated are polynomials. So a multiplication of polynomials yields a closed set of polynomials since their multiplication yields another polynomial. This can be understood through the simple addition of exponents when multiplying by said exponents, and since non-negative exponents are present, their results must also be positive (since you can't subtract exponents under these conditions and get a non-polynomial answer). "

OpenStudy (boldjon):

k take ur time

OpenStudy (anonymous):

Okay, so basically, all of the polynomials multiplied leads to another polynomial, which means that they are closed, correct?

OpenStudy (anonymous):

@boldjon ??

OpenStudy (boldjon):

sorry, i was helping someone and yes ur right

OpenStudy (anonymous):

Alright, thanks again!

OpenStudy (boldjon):

fan me so u can call on me whenever u need help

OpenStudy (boldjon):

byeee

OpenStudy (anonymous):

@boldjon Yea I did :) bye